OpenStudy (anonymous):

solve the following equation:

OpenStudy (anonymous):

\[-dsin35= -\frac{ 1 }{ 2 }9.81 \left( \frac{ dcos35 }{ 25 } \right)^{2}\]

OpenStudy (anonymous):

Solving for \(d\), I'm assuming? \[-d\sin25^\circ=-\frac{1}{2}(9.81)\left(\frac{d\cos35^\circ}{25}\right)^2\\ d\sin25^\circ=4.905\left(\frac{d^2\cos^235^\circ}{625}\right)\\ 0.4226d=0.0053 d^2\\ d^2-79.7358d=0\\ d(d-79.7358)=0\\ ~~~~~~~~~~~~~~~~\vdots \]

OpenStudy (anonymous):

d should be 109.. how did you get rid cos^2?